OpenStudy (anonymous):

So heres a weirdish wordish problem.. XD so it says When 8 is subtracted from Ollie’s age, the result is the same as subtracting 16 from 3 times his age. How old is Ollie? So the equation would be 3x - 16 = x - 8 Then you'd subtract 16 from both sides so now its 3x = x - 24 Then subtract 1x from both sides and thats now 2x = -24 Then x = -12 ? just making sure thats right XD

OpenStudy (anonymous):

i mean add 16 to both sides

OpenStudy (misty1212):

HI!!!

OpenStudy (misty1212):

you equation is right, your solution is not!

OpenStudy (misty1212):

\[3x-16=x-8\] looks good subtract \(x\) from both sides, then ADD \(16\)

OpenStudy (anonymous):

wait... how can Ollie be -12... he wasn't even conceived yet! Ok so im wrong...

OpenStudy (anonymous):

sorry my computer was messing up XD HI

OpenStudy (anonymous):

oooh ok

OpenStudy (anonymous):

so that would be 2x = 24 and then x = 12 ?

OpenStudy (anonymous):

wait that cant be right... @misty1212

OpenStudy (anonymous):

my answer anyways...

OpenStudy (misty1212):

lets start at your equation, the one you wrote

OpenStudy (misty1212):

\[3x-16=x-8\] subtract \(x\) what do you get ?

OpenStudy (anonymous):

2x - 16 = -8

OpenStudy (anonymous):

then 2x = 8

OpenStudy (misty1212):

...

OpenStudy (anonymous):

then x = 4 ?

OpenStudy (misty1212):

\(\color\magenta\heartsuit\)

OpenStudy (anonymous):

oh cool! Thanks!

OpenStudy (misty1212):

you are welcome \(\color\magenta\heartsuit \color\magenta\heartsuit \color\magenta\heartsuit\)